OpenStudy (anonymous):

Evaluate 6x ÷ 3y - 4, if x = 20 and y = 2. :A. 16 B. 20 C. 36 D. 8

OpenStudy (zpupster):

just replace the x and y with the given values: but do the division first like so: (6(20)/3(2)) -4 = ? (120/6) -4 = ? (20)-4 =? 20-4=?

OpenStudy (anonymous):

(6*20)=120 (3*2)=6 So: 120 / 6 =20 20-4=16 16 is your answer

OpenStudy (anonymous):

Did this help